# Unification of Exceptional Points and Transmission Peak Degeneracies in a Highly Tunable Magnon-Photon Dimer

### v2

@@ IV Results
The experimental methods are summarized here, and detailed in Methods. For each value of $\Delta_{\kappa}$ ($\Delta_{f}$) for the RP (ARP) configuration sweep, we independently probe each mode in isolation (by setting $J=0$) using the microwave switch technique described in Sec. II and fit the isolated mode trace to a Lorentzian to extract cavity (YIG) parameters $f_{c(y)},\kappa_{c(y)}$ and the associated fit uncertainties $\sigma_{\text{fit}}({f_{c(y)}})$ $\sigma_{\text{fit}}({\kappa_{c(y)}})$. In general, $\sigma_{\text{fit}}$ refers to uncertainty extracted from least squares regression covariance matrix. Then, coupling is activated ($J\neq 0$), and the coupled transmission Eq. 12 is fit to the experimental data using $f_{c,y},\kappa_{c,y}$, extracting $J$ and $\sigma_{\text{fit}}(J)$ for each data point. Finally, $f_{c,y},\kappa_{c,y},J$ are fit to Eq. 23 to extract $\nu_{\pm}$ data, reported in Fig. 3iii, with error bars, $\sigma_{\text{fit}}(\nu_{\pm})$ representing uncertainty propagation from $\sigma_{\text{fit}}({f_{c,y}}),\sigma_{\text{fit}}({\kappa_{c,y}}),\sigma_{\text{fit}}(J)$ (See Methods).
The fit values of $J$ and $\kappa_{c}$ are extracted for all data points, and averaged across the entire experiment, denoted as $\mu_{\text{exp}}(J),\mu_{\text{exp}}(\kappa_{c})$. We also quantify experimental uncertainty in these parameters, $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$, based on the standard deviation of $J$ and $\kappa_{c}$ across the entire experiment, which is distinct from fit uncertainty, $\sigma_{\text{fit}}$, extracted for each data point individually. Then, $\mu_{\text{exp}}(J),\mu_{\text{exp}}(\kappa_{c}),\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$ are used to simulate a distribution of possible $\nu_{\pm}$, denoted $\mathcal{D}_{\text{exp}}(\nu_{\pm})$ (shaded regions in Fig. 3iii), with standard deviation $\sigma_{\text{exp}}(\nu_{\pm})$. Finally, $\mu_{\text{exp}}(J),\mu_{\text{exp}}(\kappa_{c})$ are used to report $\Delta_{f,\kappa}^{\text{EP}}$, $\Delta_{f,\kappa}^{\text{TPD}}$, $\Delta_{f}^{\text{NL}}$, while $\sigma_{\text{exp}}(\kappa_{c}),\sigma_{\text{exp}}(J)$ determine $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{EP}})$, $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{TPD}})$, and $\sigma_{\text{exp}}(\Delta_{f}^{\text{NL}})$.
IV.0.1 Reciprocal Phase Configuration
[…]
IV.0.2 Anti-Reciprocal Phase Configuration
In the ARP configuration, we seek to consider the case where the TPD is near the EP. To do this, we set the coupling configuration to $\phi=\pi$ and select $J=8.15(4)$ MHz with $\kappa_{c}=0.752(8)$ MHz such that $\Delta_{\kappa}\approx 0$. We sweep the YIG frequency $f_{y}$ from $5.981$ to $5.998$ GHz by varying the current through an external electromagnet, controlling $B_{z}$. As the YIG frequency is tuned, the detuning $\Delta_{f}$ approaches the ARP-TPD at $\Delta_{f}^{\text{TPD}}=16.32(7)$ MHz, when the peaks merge. Crucially, for this choice of $\kappa_{c}$ and $J$, the instability transition is also near the ARP-TPD. When $\Delta_{f}$ enters the uncertainty bounds of $\Delta_{f}^{\text{NL}}=16.29(7)$ MHz, given by $\sigma_{\text{exp}}(\Delta_{f}^{\text{NL}})=0.07$ MHz [dashed lime lines in Fig. 3f], the experimental data deviate from the model, likely due to the impending nonlinearity that arises in the unstable regime. Furthermore, the proximity of the ARP-EP, at $\Delta_{f}^{\text{EP}}=16.31(7)$ MHz to the ARP-TPD causes the Petermann noise factor at the ARP-TPD to be nearly divergent. Overall, the proximity of the ARP-TPD to the ARP-EP and instability transition introduce significant deviations between the theory model and the experimental data.
IV.0.3 Performance Metrics of TPDs
A critical implication of separating TPDs and EPs is that TPDs can be moved around in parameter space using $\kappa_{c}/J$, while the EP locations are not affected by $\kappa_{c}$ (see geometric intuition in the Supplementary Information). We have shown in Fig. 2c, f that the proximity of TPDs to EPs, controllable with $\kappa_{c}/J$, changes the Petermann noise factor at the TPD and may attribute to the quality of data in 3iii. However, although the Petermann noise factor is one component of heightened uncertainty around EPs, it is not the only consideration. Here, we analyze other robustness and performance metrics associated with TPDs, and show that while the Petermann noise factor can be suppressed by shifting a TPD away from the EP, other forms of errors that are dependent on $\kappa_{c}$ and $J$ are also relevant, and do not generally follow the same parameter dependence as the Petermann noise factor. In Fig. 4, we depict the Petermann noise factor, $\bar{K}_{2}$, as a function of $\kappa_{c}$ and $J$, as well as novel metrics for TPD performance: TPD splitting strength, TPD location uncertainty propagation, and TPD minimum resolvable peak splitting, which we define in Sec. IV.0.4-IV.0.6 sections.
Figure 4: Performance Metrics of TPDs. a, Mean diagonal Petermann noise factor ($\bar{K}_{2}$) at the RP-TPD (ARP-TPD) i(ii). b, TPD location uncertainty with experimentally extracted $\sigma_{\text{exp}}(J)=0.04$ MHz, $\sigma_{\text{exp}}({\kappa_{c}})=0.01$ MHz for the RP-TPD (i) and $\sigma_{\text{exp}}(J)=0.04$ MHz, $\sigma_{\text{exp}}({\kappa_{c}})=0.008$ MHz for the ARP-TPD (ii). c, Coefficient of $\min(\Delta_{\nu})$ at the RP-TPD (ARP-TPD) i(ii). d, RP-TPD (ARP-TPD) strength i(ii).
IV.0.4 TPD Location Uncertainty
Practical sensor operation requires accurate knowledge of $\Delta_{f}^{\text{TPD}}$ or $\Delta_{\kappa}^{\text{TPD}}$, both depending on $\kappa_{c}$ and $J$. Importantly, uncertainty in these parameters, $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$, propagates to $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{TPD}})$. Furthermore, we emphasize that at the TPD, $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{TPD}})=\sigma_{\text{exp}}(\nu_{\pm}^{\text{TPD}})$. This result is critical because $\sigma_{\text{exp}}(\nu_{\pm}^{\text{TPD}})$ cannot be calculated using standard uncertainty propagation techniques, as $\nu_{\pm}$ is non-differentiable at the TPD. This correspondence between $\sigma_{\text{exp}}(\Delta_{f,\kappa}^{\text{TPD}})$ and $\sigma_{\text{exp}}(\nu_{\pm})$ can be seen in the shaded region in Fig.3iii that encapsulates the experimental data. The first order propagation of uncertainty for the TPD locations is given by
$\sigma_{\text{exp}}({\Delta_{f(\kappa)}^{\text{TPD}}})=\sqrt{\bigg{(}\frac{\partial\Delta_{f(\kappa)}^{\text{TPD}}}{dJ}\bigg{)}^{2}\sigma_{\text{exp}}^{2}(J)+\bigg{(}\frac{\partial\Delta_{f(\kappa)}^{\text{TPD}}}{d\kappa_{c}}\bigg{)}^{2}\sigma_{\text{exp}}^{2}(\kappa_{c})}.$ (24)
We stress that Eq. 24 is itself dependent on the values of $\kappa_{c}$ and $J$ (see Supplementary Information). Thus, with constant $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$, uncertainty propagation to $\sigma_{\text{exp}}({\Delta_{f,\kappa}^{\text{TPD}}})$ could be minimized by judiciously selecting $J$ and $\kappa_{c}$. Panels depicting TPD location uncertainty propagation for the RP (ARP) configurations are included in Fig. 4b.i(ii).
IV.0.5 TPD Splitting Strength
@@
The concept of splitting strength was initially introduced for EP-based sensors in Ref. [64] to quantify the pre-factor of the square root splitting with respect to the perturbation being measured. Specifically, near an EP, the splitting of imaginary eigenvalues follows $|\text{Im}(\Delta_{\lambda})|\approx 2\sqrt{J}\sqrt{\epsilon}$, where $\epsilon$ denotes the perturbation strength and $2\sqrt{J}$ is the splitting strength. Recent work has extended this concept to the ARP-TPD [25]. Here, we extend this concept to other TPDs by analyzing the coefficient of the square root term of the Puiseux series expansion of $\Delta_{\nu}$ around the TPD, with respect to $\epsilon$, a perturbation which is a function of $\Delta_{\kappa}$ ($\Delta_{f}$) for the RP (ARP) configuration. The expansion yields the best approximation of $\Delta_{\nu}(\epsilon)\approx a\sqrt{\epsilon}+\mathcal{O}(\epsilon^{3/2})$ in the proximity of each TPD for sufficiently small $\epsilon$, where $a$ is the strength. We find that this strength coefficient can be expressed as
@@
$\Delta_{\nu}^{\phi=0}(\epsilon(\Delta_{\kappa}))$ $\approx\sqrt{2}\left(8J^{2}-\kappa_{c}^{2}\right)^{1/4}\sqrt{\epsilon(\Delta_{\kappa})},$ (25)
@@
$\Delta_{\nu}^{\phi=\pi}(\epsilon(\Delta_{f}))$ $\approx\underbrace{\sqrt{2}\left(4J^{2}+\kappa_{c}^{2}\right)^{1/4}}_{\text{strength coefficient}}\sqrt{\epsilon(\Delta_{f})}.$ (26)
The expression given in Eq. 25 shows that the splitting vanishes for RP configurations with $\kappa_{c}\geq\sqrt{8}J$. Thus, the invalid domain where there is no splitting is excluded from Fig. 4a-d. Panels depicting TPD strength are shown in Fig. 4d.
IV.0.6 Minimum Resolvable Peak Splitting
The observant reader will notice that the split peaks of the experimental data in Fig. 3iii have an abrupt jump near the TPD, indicating that $\Delta_{\nu}$ fails to smoothly approach zero at the TPD itself. We refer to this effect as $\min(\Delta_{\nu})\neq 0$, which is arguably the most destructive model error for sensor performance. To understand $\min(\Delta_{\nu})$ for true TPDs versus imperfect TPDs, we begin by recalling that a true TPD is a triple-degenerate root in the cubic equation defined in Eq. 13 and requires crossing $\text{Disc}=0$ along the path $q=0$, which is $\Delta_{f}=0$ ($\Delta_{\kappa}=0$) for the RP-TPD (ARP-TPD). Satisfying this stringent requirement causes the peak splitting $\Delta_{\nu}$ to be precisely 00 at the TPD, expressed as $\min(\Delta_{\nu})=0$. However, in any experimental apparatus, this is a challenging task. Here, we consider what happens to $\min(\Delta_{\nu})$ in the case of deviation from the $q=0$ path, a manifestation of TPD parameter hypersensitivity.
If the $\text{Disc}=0$ contour is crossed with a path in parameter space where $q\neq 0$, peak splitting will still occur, but not with a continuous square root profile. Instead, the splitting is zero until it jumps to $\min(\Delta_{\nu})\neq 0$, increasing from there with a profile much closer to linear splitting than square root splitting. Mathematically, crossing $\text{Disc}=0$ with $q\neq 0$ is only a double degeneracy, with two repeated roots, denoted $r$ while one root remains a simple, separated root $s$. We quantify the distance between these roots $\min(\Delta_{\nu})=|r-s|$ as the minimum resolvable splitting, which equates to zero for a triple degeneracy (a true TPD). For the depressed cubic defined in Eq. 13, the same François Viète returns with formulas that relate the coefficients of the polynomial to sums and products of its roots. We use the Viète equations to relate $r,s,p,q$, solving $2r+s=0,r^{2}+2rs=p,r^{2}s=-q$. We find that the minimum splitting can be expressed in terms of $q$ as
$\min(\Delta_{\nu})=\frac{3}{2^{1/3}}|q|^{1/3}.$ (27)
Thus, the inability to resolve square root splitting with respect to the sensing parameter grows with the cube root of $|q|$. To quantify the notion of strength of this error, $|q|$ is expressed as a function of $\Delta_{f}$ ($\Delta_{\kappa}$) for the RP-TPD (ARP-TPD), then approximated as the first (cubic order) term of the Puiseux expansion of $|q|^{1/3}$ for the two TPDs, isolating the response as a function of $|\Delta_{f}^{1/3}|$ or $|\Delta_{\kappa}^{1/3}|$. This coefficient is expressed as
$\min(\Delta_{\nu})^{\phi=0}$ $\approx\frac{3\cdot 2^{2/3}\,\left|4J^{2}-\kappa_{c}^{2}\right|^{1/3}}{4}\,|\Delta_{f}|^{1/3},$ (28)
$\min(\Delta_{\nu})^{\phi=\pi}$ $\approx\underbrace{\frac{3\,\kappa_{c}^{1/3}\,\left(4J^{2}+\kappa_{c}^{2}\right)^{1/6}}{2}}_{\min(\Delta_{\nu})\text{ coefficient}}\,|\Delta_{\kappa}|^{1/3}.$ (29)
Panels depicting the $\min(\Delta_{\nu})$ coefficient for the two TPDs are included in Fig. 4c, g.
@@ IV.0.6 Minimum Resolvable Peak Splitting
The performance metrics introduced in this section can be used as a guide to optimize TPD-based sensors. For example, the RP-TPD can be placed along a line where $\bar{K}_{2}=1$, minimizing the Petermann noise factor and $\min(\Delta_{\nu})$ coefficient while increasing model uncertainty propagation. For the ARP-TPD, $\min(\Delta_{\nu})$ and TPD strength can be increased by moving the ARP-TPD closer to the ARP-EP, which drastically increases the Petermann noise factor. Ultimately, the design of TPD-based sensors must balance competing metrics based on the constraints and goals of the specific experimental platform.
V Discussion and Outlook
This work presents a highly tunable experimental platform and a unified theoretical framework for systematically investigating EPs and TPDs in coupled oscillator systems. A synthetic gauge field is engineered through the complex coupling term $Je^{i\phi}$, effectively unifying previously distinct two-dimensional configurations of EPs and TPDs into an exceptional surface. This synthetic gauge field enables continuous tuning between RP and ARP coupling regimes, and we report the distinct properties of the RP-EP, RP-TPD, ARP-EP, and ARP-TPD, highlighting their differences through experimental and theoretical analysis.

### v3

@@ III Experimental Results
The experimental methods are summarized here, and detailed in Methods. For each value of $\phi$, we record $J$ and use it to convert measured frequencies (such as $\Delta_{f}$) to nondimensionalized parameters (such as $\tilde{\Delta}_{f}$). Sweeps are initialized by setting $\tilde{\kappa}_{c}$ and specifying a $\tilde{\Delta}_{f}$ or $\tilde{\Delta}_{\kappa}$ range. A control system is used to follow a trajectory in $\tilde{\Delta}_{f}$-$\tilde{\Delta}_{\kappa}$ space (Methods). For each step of a sweep, we independently probe each mode in isolation (by setting $J=0$) and fit the isolated mode trace to a Lorentzian, extracting cavity (YIG) parameters $f_{c(y)},\kappa_{c(y)}$ and $\sigma(f_{c(y)}),\sigma(\kappa_{c(y)})$ to obtain $\tilde{\Delta}_{f},\tilde{\Delta}_{\kappa},\sigma(\tilde{\Delta}_{f}),\sigma(\tilde{\Delta}_{\kappa})$. Then, coupling is enabled $(J\neq 0)$, and Eq. 2 is fit to the experimental data using $\tilde{\Delta}_{\kappa}$ and $\tilde{\Delta}_{f}$. All parameters are fit to Eq. 12 to extract $\tilde{\nu}_{\pm}$ data, reported in Fig. 3iii, with error bars representing uncertainty propagation from $\sigma(\tilde{\Delta}_{\kappa}),\sigma(\tilde{\Delta}_{f})$.
For $\phi=0$, we calibrate $J=1.05$ MHz (Methods), and $\tilde{\Delta}_{\kappa}$ is swept while $\tilde{\Delta}_{f}=0$, with $\tilde{\Delta}_{\kappa}^{\text{EP}}=\pm 2$, $\tilde{\Delta}_{\kappa}^{\text{TPD}}=\frac{\tilde{\kappa}_{c}}{2}\pm\frac{\sqrt{8-\tilde{\kappa}_{c}^{2}}}{2}$, $\tilde{\Delta}_{f}^{\text{EP, TPD}}=0$, provided $\tilde{\kappa}_{c}^{2}-8\leq 0$. Figure 3a-b depicts the EP at $\tilde{\Delta}_{\kappa}^{\text{EP}}=-2$, with TPDs at $\tilde{\Delta}_{\kappa}^{\text{TPD}}=-1.04(1)$ for $\tilde{\kappa}_{c}=0.67(1)$ (Fig. 3a) and $\tilde{\Delta}_{\kappa}^{\text{TPD}}=-0.04(4)$ for $\tilde{\kappa}_{c}=1.96(1)$ (Fig. 3b). The peak locations, $\tilde{\nu}_{\pm}$, do not split at the EP, and even after they split at the TPD, they remain bounded by the imaginary eigenvalues (Fig. 3a.i, b.i). The instability transition occurs at $\tilde{\Delta}_{\kappa}^{\text{NL}}=\tilde{\kappa}_{c}$ for $\kappa_{c}\leq 2$, or $\tilde{\Delta}_{\kappa}^{\text{NL}}=\frac{\tilde{\kappa}_{c}^{2}+4}{2\tilde{\kappa}_{c}}$ for $\tilde{\kappa}_{c}>2$. Figure 3a.i-iii shows $\tilde{\Delta}_{\kappa}^{\text{NL}}=0.67(1)$.
For $\phi=\pi$, we calibrate $J=1.13$ MHz, and $\tilde{\Delta}_{f}$ is swept while $\tilde{\Delta}_{\kappa}=0$, with $\tilde{\Delta}_{f}^{\text{EP}}=\pm 2$, $\tilde{\Delta}_{f}^{\text{TPD}}=\pm\sqrt{\tilde{\kappa}_{c}^{2}+4}$, $\tilde{\Delta}_{f}^{\text{NL}}=\pm\sqrt{4-\kappa_{c}^{2}}$, $\tilde{\Delta}_{\kappa}^{\text{EP, TPD, NL}}=0$. Figure 3c-d depicts the EP at $\tilde{\Delta}_{f}^{\text{EP}}=+2$, with TPDs and instability transitions at $\tilde{\Delta}_{f}^{\text{TPD}}=2.167(4),\tilde{\Delta}_{f}^{\text{NL}}=1.817(4)$ for $\tilde{\kappa}_{c}=0.83(1)$ (Fig. 3c) and $\tilde{\Delta}_{f}^{\text{TPD}}=2.598(7),\tilde{\Delta}_{f}^{\text{NL}}=1.117(7)$ for $\tilde{\kappa}_{c}=1.66(1)$ (Fig. 3d).
For intermediate values of $\phi$, including $\phi=\pi/2$, $\tilde{\Delta}_{f}$ and $\tilde{\Delta}_{\kappa}$ must be simultaneously swept along a hyperbolic trajectory defined by $\tilde{\Delta}_{\kappa}\tilde{\Delta}_{f}=2\sin(\phi)$ ($\tilde{q}=0$ in Eq. 11). This trajectory crosses both the EP and the TPD, and is achieved using a parameter control system (Methods), with $(\tilde{\Delta}_{\kappa}^{\text{EP}},\tilde{\Delta}_{f}^{\text{EP}})=(\pm\sqrt{2},\pm\sqrt{2})$. Experimentally, we calibrate $\phi=\pi/2,J=0.95$ MHz, and $\tilde{\Delta}_{\kappa}$ is chosen as the independent variable, with the control system automatically adjusting $\tilde{\Delta}_{f}$ at each step to stay on the hyperbolic trajectory. Figure 3e-f depicts the EP at $(\tilde{\Delta}_{\kappa}^{\text{EP}},\tilde{\Delta}_{f}^{\text{EP}})=(-\sqrt{2},-\sqrt{2})$, projected onto the $\tilde{\Delta}_{\kappa}$ axis. The TPD locations are found from the roots of a quartic equation, with coefficients dependent on $\phi,\tilde{\kappa}_{c}$, and cannot be written in closed form (see Supplementary Information for details). Figure 3e depicts the TPD at $(\tilde{\Delta}_{\kappa}^{\text{TPD}},\tilde{\Delta}_{f}^{\text{TPD}})=(-0.860(3),-2.327(9))$ for $\tilde{\kappa}_{c}=1.30(1)$, Fig. 3f has $(\tilde{\Delta}_{\kappa}^{\text{TPD}},\tilde{\Delta}_{f}^{\text{TPD}})=(-0.657(4),-3.04(2))$ for $\tilde{\kappa}_{c}=2.32(3)$ (both projected onto the $\tilde{\Delta}_{\kappa}$ axis). As shown in Fig. 2c.iii-iv, the instability transition does not occur within a meaningful operating regime for the $\phi=\pi/2$ primary configuration.
III.1 Figures of Merit of Transmission Peak Degeneracies
A critical implication of separating EPs and TPDs is that TPDs can be moved around in parameter space using $\tilde{\kappa}_{c}$, without changing the location of EPs. We have shown in Fig. 2iii-iv that the proximity of TPDs to EPs changes the Petermann noise factor at the TPD. This relationship is illustrated for a range of $\tilde{\kappa}_{c}$ values in Fig. 4iv-v, with Fig.4iv depicting the euclidean distance in the $\tilde{\Delta}_{\kappa}$-$\tilde{\Delta}_{f}$ plane between the primary TPD and the nearest EP, and Fig. 4v depicting the Petermann noise factor at the TPD. Circles (triangles) in Fig. 4b represent $\tilde{\kappa}_{c}$ values chosen for Fig. 3a,c,e(b,d,f). At $\phi=0,\tilde{\kappa}_{c}=2$, a TPD occurs with minimal $\bar{K}_{2}$ and thus an orthogonal eigenbasis. Here we report other useful figures of merit for TPDs, and show that while $\bar{K}_{2}$ can be reduced by shifting a TPD away from the EP, other relevant characteristics that depend on $\phi$ and $\tilde{\kappa}_{c}$ can be used to describe the performance of TPDs.
@@ III.1 Figures of Merit of Transmission Peak Degeneracies
The concept of splitting strength was initially introduced for EP-based sensors in Ref. [51] to quantify the pre-factor of the square root splitting with respect to the perturbation being measured. Specifically, near an EP, the splitting of imaginary eigenvalues follows $|\text{Im}(\tilde{\Delta}_{\lambda})|\approx 2\sqrt{\tilde{\epsilon}}$, where $\tilde{\epsilon}=\epsilon/J$ for a perturbation $\epsilon$ in either $\Delta_{\kappa}$ or $\Delta_{f}$. For $\phi=\pi/2$, the EP strength is still $2$ as long as $\tilde{\epsilon}$ is a perturbation along the hyperbolic trajectory $\tilde{q}=0$. Recent work has reported the splitting strength of the $\phi=\pi$ TPD [5]. Here, we extend this concept to all TPDs by analyzing the Puiseux series expansion of $\tilde{\Delta}_{\nu}$ around the TPD, with respect to $\tilde{\epsilon}$, and find
@@
$\tilde{\Delta}_{\nu}^{\phi=0}(\tilde{\epsilon}(\tilde{\Delta}_{\kappa}))$ $\approx\sqrt{2}\left(8-\tilde{\kappa}_{c}^{2}\right)^{1/4}\sqrt{\tilde{\epsilon}(\tilde{\Delta}_{\kappa})},$ (14)
@@
$\tilde{\Delta}_{\nu}^{\phi=\pi}(\tilde{\epsilon}(\tilde{\Delta}_{f}))$ $\approx\underbrace{\sqrt{2}\left(\tilde{\kappa}_{c}^{2}+4\right)^{1/4}}_{\text{strength coefficient, }\tilde{a}}\sqrt{\tilde{\epsilon}(\tilde{\Delta}_{f})}.$ (15)
For $\phi=\pi/2$, the strength cannot be written in closed form, but is plotted in Fig. 4b.ii along with Eq. 14 and Eq. 15. We stress that for the $\phi=\pi/2$ TPD, the strength coefficient is normalized based on a perturbation exactly along the hyperbolic path $\tilde{q}=0$. If strength is projected onto either the $\tilde{\Delta}_{\kappa}$ or $\tilde{\Delta}_{f}$ axis, the splitting is significantly stronger than for the $\phi=0,\pi$ TPD.
Parameter hypersensitivity exists as a challenge for EPs and was recently explored for the $\phi=0$ TPD [19]. Here, we report one quantitative measure of the effect of hypersensitivity on TPDs. The split peaks of the experimental data in Fig. 4a have an abrupt jump near the TPD, indicating that $\tilde{\Delta}_{\nu}$ fails to smoothly approach zero at the TPD itself. If the TPD is crossed with a path in parameter space where $\tilde{q}\neq 0$, due to fabrication or operational errors, peak splitting will still occur, but not with a continuous square root profile. Instead, the splitting is zero until it jumps to $\min(\tilde{\Delta}_{\nu})\neq 0$, increasing from there. Mathematically, we find that the minimum splitting can be expressed in terms of $\tilde{q}$ as
$\min(\tilde{\Delta}_{\nu})=\frac{3}{2^{1/3}}|\tilde{q}|^{1/3},$ (16)
with a derivation involving the geometry of the roots of Eq. 9 in the Supplementary Information. $\min(\tilde{\Delta}_{\nu})=\tilde{a}\sqrt{\tilde{\epsilon}_{\text{min}}}$ bounds the minimum perturbation that can be resolved, limiting the precision of the TPD. The exact value of $\min(\tilde{\Delta}_{\nu})$ is based on how far off the $\tilde{q}=0$ path an experiment strays. For the $\phi=0$ ($\phi=\pi$) TPD, this occurs when $\tilde{\Delta}_{f}\neq 0$ ($\tilde{\Delta}_{\kappa}\neq 0$). Our calculations use the convergence threshold of the control system to calculate the maximum $\tilde{q}$ deviation (Methods). $\min(\tilde{\Delta}_{\nu})$ is plotted in Fig. 4b.iii, illustrating its dependence on $\phi$ and $\tilde{\kappa}_{c}$.
The promise of EP-based sensors revolves around the potential to attain an infinite derivative response exactly at the onset of square root splitting. For TPDs, splitting strength and minimum splitting combine into a more realistic picture for the maximum possible derivative. Around the TPD, $\tilde{\Delta}_{\nu}\approx\tilde{a}\sqrt{\tilde{\epsilon}}$, where $\tilde{a}$ is the normalized strength of the TPD, with derivative ${\tilde{a}}/{2\sqrt{\tilde{\epsilon}}}$. The maximum value of this derivative is based on the minimum resolvable $\tilde{\epsilon}$, with $\text{Max Derivative}=\tilde{a}^{2}/(2\min(\tilde{\Delta}_{\nu})$ plotted in Fig. 4b.i.
@@
The figures of merit introduced in this section can be used as a guide to optimize TPD-based sensors. For example, the $\phi=0$ TPD should be used with $\tilde{\kappa}_{c}=2$, minimizing $\bar{K}_{2}$ and $\min(\tilde{\Delta}_{\nu})$ at the cost of some splitting strength. For the $\phi=\pi$ TPD, response can be maximized by moving the TPD closer to the EP, which inadvertently drastically increases the Petermann noise factor. Ultimately, the design of TPD-based sensors must balance competing metrics based on the constraints and goals of the specific experimental platform.
Figure 4: Figures of Merit for TPDs. a-c, Experimental data from Fig. 3iii, recast as $\tilde{\Delta}_{\nu}=\tilde{\nu}_{+}-\tilde{\nu}_{-}$ for representative small $\tilde{\kappa}_{c}$ (circle) and large $\tilde{\kappa}_{c}$ (triangle) directly compared at $\phi=0$ (a), $\phi=\pi$ (b), $\phi=\pi/2$ (c). Horizontal lines depict theoretical $\min(\tilde{\Delta}_{\nu})$ quantified using uncertainties in the experimental apparatus. Data illustrate that the pre-factor of the square root splitting, as well as the minimum resolvable splitting, depend on $\tilde{\kappa}_{c}$. d depicts figures of merit for TPDs, including i maximum derivative, ii TPD splitting strength, iii minimum resolvable splitting, iv, euclidean distance from TPD to nearest EP, and v Petermann noise factor ($\bar{K}_{2}$) at the TPD.
IV Conclusion and Outlook